15. Wh en computing tax depreciation in the first year for new equipment that has a recovery period of less than 20 years, you a. must take 50% bonus depreciation unless you elect not to do so. b, cannot take 50% bonus deprec iation unless you elect to do so. c. can or cannot take 50% bonus depreciation as you wish; no election is required. d must take 50% bonus depreciation unless you elect not to do so, but must wait to take normal Table 1 depreciation until Year 2. 16. Under GAAP, annual depreciation for a building can be allocated . . a. entirely to Depreciation Expense, entirely to Inventory- Work-In-Process OH or partly to both, depending on how the building is used. b. only to Depreciation Expense. c. only to Inventory-Work-In-Process OH d. entirely to either Depreciation Expense or Inventory Work-In-Process OH, but not allocated partly to both. 17. To depreciate an asset under the double-declining balance method multiply... a. the depreciable base by each years depreciation rate b. the depreciation rate by each years beginning book value. c. the acquisition cost by each years depreciation rate d. the cost basis by each years depreciation rate. Under straight-line depreciation, the annual depreciation rate is computed by... 18. a. dividing 100% by the estimated life. b. dividing the numeral 1.00 by the estimated life. c. either a or b. d. none of the above 19. Under sum-of-the-years-digits depreciation... a. the book value remains the same each year. b. the depreciation rate changes each year. c. the denominator of the SYD fraction changes each year. d. all of the above.
Answer #1

15.) Answer - A

The tax payers are required to take additional bonus depreciation of 50% in first year on new equipment that has a recovery period of less than 20 years unless you elect as not to do so (if you does not take the additional bonus depreciation)

16.) Answer - A

Allocated entirely to depreciation expense if some part of the building is used for other than manufacturing purpose and  if some part of the building is used for the manufacturing purpose then it is allocated entirely to inventory - Work-In-Process OH. It is depending on how the building is used.

17.) Answer - B

In this method, "double" means 200% of straight line rate of depreciation and "declining balance" means assets book value or carrying value at the beginning of the year. so, each year beginning book value or carrying value is calculated with 2x rate ( double than straight line depreciation rate).

18.) Answer - C

As the question is about annual depreciation rate computation in straight line depreciation, it can be calculated either by dividing 100% by estimated life or by dividing the numeral 1.00 by the estimated life of the asset.

example - life of the asset is 5 years under straight line method. so the annual depreciation rate can be calculated as

1.) 100% / 5 years =0.20 *100 = 20 % depreciation rate per year

or

2.) 1/5 of the estimated life of asset = 0.20*100 = 20% depreciation rate per year.

19.) Answer - B

formula : Asset Depreciable Amount * (remaining useful life / sum of years digits)

so, here the rate of depreciation changes each year as the remaining useful life decrease and the sum of years digits (denominator) is same for every year. (the book value of asset falls each year with increase in balance in accumulated depreciation account.

1